OpenStudy (anonymous):

A footbridge is in the shape of an arc of a circle. The bridge is 6 ft tall and 25 ft wide. What is the radius of the circle that contains the bridge?

OpenStudy (anonymous):

wouldnt it be half of 25?

OpenStudy (anonymous):

maybe, but then whats the 6 for?

OpenStudy (anonymous):

thats what i couldnt figure out

OpenStudy (anonymous):

haha it's confusing.

OpenStudy (anonymous):

i'll keep trying ha

OpenStudy (anonymous):

i really dont know though sorry!

OpenStudy (amistre64):

i do :)

OpenStudy (anonymous):

it okaye michelle(: haha and can you help then.? :P

OpenStudy (amistre64):

w^2 ----- + rise/2 its how i made a thousand arches rise/8

OpenStudy (amistre64):

25^2/6 = n n/8 = m m+ 6/2 = radius

OpenStudy (amistre64):

abt 16.02

OpenStudy (amistre64):

we would have walkthrus that the prints needed arches so high by so wide; i figured id find a formula for determing how far to put my pivot point to attc my chalkline or tape to so I could make the required arch

OpenStudy (amistre64):

width^2 ------ + rise/2 = radius of circle rise/8